OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal!!

OpenStudy (abbycross167):

Please help me to simplify this rational expression - (6x) (2y+2) ------ * ------- (y+1) (3)

jhonyy9 (jhonyy9):

so i m here

jhonyy9 (jhonyy9):

rewrite these terms 3*2x 2(y+1) ------ * ------- = y+1 3 what can you simplifie cross there ?

jhonyy9 (jhonyy9):

do you see it ? 3 by 3 and what again ?

jhonyy9 (jhonyy9):

and y+1 by y+1 right so and what will get ?

OpenStudy (abbycross167):

@jhonyy9 do you want me to cross multiply?

OpenStudy (abbycross167):

If so then - 3*3= 9 And I think y+1*y+1=y+1? Right?

jhonyy9 (jhonyy9):

no multiplie cross simplifie

jhonyy9 (jhonyy9):

3 by 3 and y+1 by y+1 and so what will reamin there again ?

jhonyy9 (jhonyy9):

remain